Find out the time at which the potential across the capacitor is equal to that across the
resistor, when switch S is closed at t=0, the capacitor is charged through the resistor.
Consider the value of capacitance of a capacitor= 12mF and resistance of resistor is 25ohm.
1
Expert's answer
2020-08-13T12:24:53-0400

The voltage across a capacitor during charging can be describe with the following formula:


"U_C = U_0(1 - e^{-\frac{t}{RC}})"

where "U_0" is the voltage of the battery, "t" is time, "C = 12\times 10^{-3}F" is the capacitance of the capacitor and "R = 25\Omega" is the resistance of the resistor.

On the other hand, the voltage across the resistor is:


"U_R = U_0 - U_C = U_0 - U_0(1 - e^{-\frac{t}{RC}}) = U_0e^{-\frac{t}{RC}}"

The required condition is:


"U_C = U_R\\"

"U_0(1 - e^{-\frac{t}{RC}}) = U_0e^{-\frac{t}{RC}}\\
e^{-\frac{t}{RC}} = \dfrac12"

Expressing the time from the last equation, get:


"t = -RC\ln\dfrac12 = -25\cdot 12\cdot 10^{-3}\ln \dfrac12\approx 0.21s"

Answer. 0.21 s.